Grade 11Mechanics

A point moves rectinearly. Its position x at time t is given by x^2=t^2+1. Its acceleration at time t is(A) 1/x^3(B) 1/x-1/x^2(C)-t/x^2(D)none of these

Kashish
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
(A) option.
Differentiate x w.r.t t once,
dx/dt= t/√(t^2+1)
=>d2x/dt2 = a = (t2-t2+1)/(t2+1)3/2
=>a=1/x3.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)